Lilian9667:

2(a+3)=-4 A.) -5 B.) -2 C.) 3 D.) 4

Lilian9667:

There that's better

supie:

Yeah so the first step would be to distribute so we already have \(2(a+3)=−4\) \((2)(a)+(2)(3)=−4\) Which equals what?

supie:

Basically all you're doing is solving the multiplication (2)(a)+(2)(3)=−4 ^ ^ 2xa 2x3

Lilian9667:

Is this adding

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 Is this adding \(\color{#0cbb34}{\text{End of Quote}}\) multiplication

Lilian9667:

or Multiplication

supie:

Multiplication

Lilian9667:

ohh

Lilian9667:

lol

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Yeah so the first step would be to distribute so we already have \(2(a+3)=−4\) \((2)(a)+(2)(3)=−4\) Which equals what? \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Basically all you're doing is solving the multiplication (2)(a)+(2)(3)=−4 ^ ^ 2xa 2x3 \(\color{#0cbb34}{\text{End of Quote}}\)

supie:

Confused...?

Lilian9667:

here give me a moment hmm?

supie:

ight

Lilian9667:

12? and where's the post to be by the 2x a<

Lilian9667:

I mean whats the a post to be by the 2

supie:

What?

Lilian9667:

it says 2xa <

Lilian9667:

a <<<<<

supie:

All you're doing is multiplying 2 and \(a\) and multiplying 2 and 3

supie:

Confused...again....?

Lilian9667:

Yah

supie:

So we're multiplying 2 and 3 which is 6 and 2 and a which is 2a so that would leave is with \(2a+6=−4\)

Lilian9667:

6

Lilian9667:

oh so it's 4

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 oh so it's 4 \(\color{#0cbb34}{\text{End of Quote}}\) what's 4?

Lilian9667:

Haha Nevermind

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 Haha Nevermind \(\color{#0cbb34}{\text{End of Quote}}\) lol ok

Lilian9667:

I thought that was the answer

supie:

So now do you know the next step what we have to do to \(2a+6=−4\)?

Lilian9667:

no

supie:

Ok so we have to change the `+` to a `-` and we have to take the `6` and subtract it from both sides of the equation the reason we `subtract` is because we changed the `+` to a `–`

supie:

So we changed the plus to a minus no we're subtracting 6 from both sides of the equation \(2a+6−6=−4−6\) so that equals what?

supie:

Still confused....?

Lilian9667:

Oh sorry just asking my mom about this question

supie:

Oh ok

Lilian9667:

yah She said to go to khan academy? I guess

Lilian9667:

CUS THIS IS HARD DON"T KNOW HOW YOU CAN DO THIS

Lilian9667:

Huh

SkilerLuke:



Lilian9667:

-5

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 -5 \(\color{#0cbb34}{\text{End of Quote}}\) yeah good job

SkilerLuke:

yes

supie:

You can close this now @Lilian9667

Lilian9667:

Ok